OpenStudy (legends):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie.

OpenStudy (legends):

c(x)=1.50 + 1.25= x c(x)=1.50x + 1.25 c(x)=2.75 c(x)= (1.50 + 1.25)x

OpenStudy (legends):

@iGreen @rvc @TheSmartOne

OpenStudy (legends):

The first option is wrong sorry its like this c(x)=1.50 + 1.25x

OpenStudy (baby_bear69):

no matter what she'll be paying $1.50 since that's the flat fee. So it's $1.50 plus 1.25 x how many nights she keeps it. So if x is the amount of days she is keeping the movie, you're going to multiply that by $1.25. So c(x)= 1.50 + 1.25x

OpenStudy (legends):

Ooooohhh okay I got it thanks Pineapple :D

OpenStudy (baby_bear69):

Anytime, Mango xD